OpenStudy (anonymous):

PLEASE HELP MEDALS INCLUDED 4/11 * 10/8 a. 4/11 b.5/11 c. 6/11

OpenStudy (anonymous):

first you gotta simplify 10/8

OpenStudy (anonymous):

5/4

OpenStudy (anonymous):

Don't report him, but know that we usually want the questoneer to answer problems. Just assist them on getting the question right.

Mehek (mehek14):

easiest way to do this multiply (4/11) by 8 and (10/8) by 11 then subtract and simplify

OpenStudy (anonymous):

this whole thing is wrong because there are no negative options, and if you look at it 4/11-10/8 =something negative

OpenStudy (anonymous):

um no wonder I've been having a hard time doing long division and stressing tf out

OpenStudy (anonymous):

this is about helping them figure the answer out not just giving them the answer it's against the code of conduct, she's not learning anything

Mehek (mehek14):

hmmm lets see \(\dfrac{4}{11}=\dfrac{32}{88}\) \(\dfrac{10}{8}=\dfrac{110}{88}\) \(\dfrac{32-110}{88}=\dfrac{-78}{88}=\dfrac{-39}{44}\) @sheyennepowell_ you sure the question is written correctly?

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

i don't think so

OpenStudy (anonymous):

oops its not a minus its multiplication

Mehek (mehek14):

o.o lol k so 4*10 and 8*11

OpenStudy (anonymous):

hahahahahahaha sorry

Mehek (mehek14):

\(\dfrac{4*10}{11*8}=?\)

OpenStudy (anonymous):

https://www.khanacademy.org/math/arithmetic/fractions/multiplying_fractions/v/multiplying-fractions watch this short video and you'll understand, if you get it right. A medal would be greatly appreciated

OpenStudy (anonymous):

thats what i got dont know what to do